OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

for y-intercept, x coordinate is zero. so substitute x=0 and get y-intercept y=-2(0)-3 = -3 so y-intercept is (0, -3) for x-intercept, y coordinate is zero. so substitute y=0 and get x-intercept 0=-2x-3 => 2x = -3 => x = -3/2 so x-intercept is (-3/2, 0)

OpenStudy (anonymous):

oooo thanks! i understand it now..so those are my answers right

OpenStudy (anonymous):

Yes the last lines give the coordinates of y-intercept and x-intercept

OpenStudy (anonymous):

ooh yeah thanks sooo much